OpenStudy (anonymous):

cos 2x - cos x = 0. Find all solutions in interval [0,2pi]. I literally have no idea how to do this so a link to a youtube video or an explanation would be awesome

OpenStudy (anonymous):

\[\cos2x=2\cos ^{2}x-1\] \[(2cosx+1)(cosx-1)=0\] \[\cos x=-0.5,1\] between 0 and 2pi, therefore the solutions are x=0, 2pi, 2pi/3, 4pi/3